4log(base 3)(2x+1) -7 = 1
solve for x

Answer by ewatrrr(24785)

 
Hi

4log3 2x+1) -7 = 1
4log3(2x+1)= 8
log3(2x+1)= 2
3^2 = 2x+1
9 = 2x+1
8/2 = x = 4